OpenStudy (anonymous):

how do I graph x - 1 < -11 or x + 12 > 14

terenzreignz (terenzreignz):

First, you need to simplify them both...

OpenStudy (anonymous):

work out x by going x-1=0 and x+12=0 and then solve for y by substituting in 0 for x (e.g. x-1=y, 0-1=y) etc

OpenStudy (anonymous):

for first eq x - 1 < -11 x<-10 for the second x+12>14 or x>2 that means u have x<-10 or x>2 on the number line first locate -10 and 2 mark the regions towards left of -10 (starting from -10 but excluding -10) similarly mark the regions towards right of 2 (starting from 2 but excluding 2)